# From Perception to Reasoning: Deep Thinking Empowers Multimodal Large Language Models

### v1

@@ 3.3.1 CoT Prompting
Compared to directly generating an answer, CoT prompting, through few-shot example learning or zero-shot prompts, induces the model to spontaneously form a reasoning chain, improving its performance on multi-step reasoning tasks. Wei et al.(wei2022chain) first introduced CoT prompting method in LLMs, improving model performance on arithmetic, commonsense, and logical reasoning tasks. Kojima et al.(kojima2022large) further proposed the Zero-shot CoT prompting method, which only requires adding the cognitive trigger instruction ”Let’s think step by step” to the prompt to activate the model’s internal serialization and logical inference pathways, successfully guiding its reasoning pattern.
In multimodal scenarios, the core challenge for CoT prompting lies in providing a guiding framework to synergistically process and transform information from different modalities to bridge the representation gap. The BBA method(zhao2024bba) guides a vision-language model to create separate reasoning chains for visual representations and Domain-Specific Language (DSL) representations, and then integrates the information from different modalities by aligning inconsistencies; this decoupling-alignment strategy reduces cognitive load and problem complexity. CCoT(mitra2024compositional) first utilizes a Multimodal Large Language Model to generate a scene graph, which is then embedded into the prompt to generate an answer. The IoT prompting method by Zhou et al.(zhou2024image) guides the model to automatically design visual information extraction operations based on the image and question, acquiring visual auxiliary information for each reasoning step. This transforms the model from a passive receiver of information into an active agent that acquires information, making the reasoning results more credible. Furthermore, prompting methods such as DDCoT(zheng2023ddcot), MI-CoT(anand2024mm), and ICoT(gao2025interleaved) incorporate visual cues in the prompts, compelling the model to trace back and verify visual evidence at each step of language generation. This achieves a continuous grounding of linguistic symbols with visual information, effectively suppressing model hallucinations.s.
@@ 3.3.2 Search Strategies
Autoregressive generation models, during inference, follow a greedy strategy, meaning they select the locally optimal next token at each step. This decision-making mechanism is highly susceptible to compounding errors, where minor early deviations are continuously amplified in subsequent steps, ultimately deviating from the correct reasoning trajectory. Search strategies in CoT reasoning aim to systematically overcome this limitation. The core idea is to expand the single, fragile reasoning path into an explorable solution space at the cost of increased computation, thereby finding a more optimal reasoning path globally and improving the correctness and stability of the output. As shown in subfigure (b) of Figure 5, compared to single-path reasoning, search strategies enable the model to more comprehensively explore plausible reasoning chains by evaluating and selecting from multiple branches, enhancing its robustness and generalization capabilities in multi-step reasoning tasks. The search process for MCoT is more complex than for its textual counterpart, as it requires joint decision-making across dimensions such as modal attention, cross-modal alignment, and perceptual understanding, resulting in a higher-dimensional search space. Common search strategies include Beam Search, Monte Carlo Tree Search (MCTS), and Best-of-N (BoN) search(amini2024variational; wang2024improved).
Beam Search: As a widely used heuristic search algorithm, its core idea is to retain a fixed number (known as the ”beam width”) of the most probable candidate sequences at each step of the generation process and to expand upon them in subsequent steps. Theoretically, it is a practical compromise between greedy search and exhaustive search, mitigating the fragility of a single path through limited parallel exploration. However, its heuristic assumption may lead to missing the global optimum due to short-sightedness. In the multimodal context, to balance the granularity and complexity of the search, Xu et al.(xu2024llava) proposed a stage-level beam search method, which executes the search by leveraging the phased output characteristic of LLaVA-CoT, with its performance improving as the beam width increases.
@@
To address the reliability issue of reasoning path expansion in MCoT, AR-MCTS(dong2024progressive) integrates MCTS with an active retrieval mechanism to achieve automatic generation of step-by-step annotations. Its innovation lies in combining internal contemplation (MCTS) with external knowledge seeking (active retrieval), ensuring the search process is both logically coherent and factually grounded, thus enhancing the reliability of the exploration. Similarly, to tackle the search effectiveness and efficiency problems in multimodal search methods, Yao et al.(yao2024mulberry) proposed CoMCTS, which introduces collective learning into the tree search. It utilizes the collective knowledge of multiple models to collaboratively hypothesize, search, and identify effective reasoning paths through iterative operations, thereby smoothing out the cognitive biases of a single model and improving search coverage and the probability of finding the optimal solution.
To balance the performance and efficiency of AR-MCTS and CoMCTS, Wu et al.(wu2025boosting) proposed AStar, a multimodal automatic structured thought paradigm based on MCTS. Its contribution is to elevate the search object from low-dimensional token sequences to the level of high-dimensional, pre-defined cognitive actions, allowing for the derivation of high-level cognitive reasoning patterns with greater efficiency. Furthermore, Sun et al.(sun2025mm) proposed a simulation-based tree search algorithm. Considering that MLLMs are affected by hallucinations and struggle to produce reliable rewards, this method employs a simulation-based reward mechanism to replace the traditional one in MCTS, providing a more reliable signal for the value assessment (backpropagation step) of MCTS and enhancing the robustness of the search process.
Best-of-N Search: The basic idea of Best-of-N is that the model generates N possible outputs, and then the optimal solution is selected from them based on a preset evaluation criterion. This method increases the likelihood of obtaining a high-quality output by increasing the number of candidate solutions, making it suitable for tasks with high requirements for the quality of the generated results. The SPECULATIVE REJECTION method(sun2024fast) improves upon Best-of-N by halting the generation of low-scoring candidate answers, thereby significantly reducing the model’s demand for computational resources. The TreeBoN method(qiu2024treebon) merges a speculative tree search strategy with the Best-of-N method, achieving a balance between high output quality and efficient inference by iteratively expanding branches and pruning low-quality answers. Lin et al.(lin2025investigating) systematically studied inference-time scaling methods like Best-of-N, demonstrating that the performance of MCoT reasoning consistently surpasses that of text-only reasoning.
@@
For challenging multimodal mathematical reasoning tasks, the AtomThink framework(xiang2024atomthink) combines four search strategies and a policy reward model. It divides the search into path search (using majority voting and Best-of-N to aggregate scores and find the optimal solution) and step-wise search (gradually expanding the atomic action sampling space from an initial path, using beam search and greedy strategies to prune low-quality paths). To generate model answers with better visual understanding, Wang et al.(wang2024scaling) proposed VisVM to serve as a signal to guide the model’s search during inference. The innovation of VisVM lies in introducing a more forward-looking guidance signal, providing long-term value to avoid local optima.
Additionally, to address the problems of granularity mismatch and interference from irrelevant information in in-context learning, Zhang et al.(zhang2025booststep) proposed BoostStep, which improves reasoning accuracy through step-level alignment in in-context learning and can be seamlessly integrated with CoT and tree search algorithms. To mimic the interleaved nature of the human visual-language reasoning process, Wang et al.(wang2025visuothink) proposed the VisuoThink reasoning framework, which dynamically integrates visual and textual information. Its core is a predictive rolling search mechanism that fuses visual-textual thought, enabling timely corrections by predicting intermediate state results, thereby enhancing the model’s reasoning performance.
@@ 3.3.3 Self Refinement
In recent years, a significant body of work(dhuliawala2023chain; gou2023critic; paul2023refiner; madaan2023self; li2023making; shinn2023reflexion) has demonstrated that large models possess self-correction capabilities; that is, they can improve the quality of their answers through reflection. This paradigm transforms the model’s generation process from a linear ”think-output” pattern to an iterative ”think-output-review-refine” loop, optimizing the reasoning path through an internal feedback mechanism without updating the model’s parameters. As illustrated in subfigure (c) of Figure 5, after generating an initial response, the model reflects upon the original input and its initial answer to generate feedback. Finally, the model refines the initial response based on this feedback. Generally, self-optimization can rely on the model’s internal prompting and external information.
To address the issue that MLLMs struggle to learn effectively from erroneous reasoning, R3V(cheng2024vision) introduces self-correction and self-selection, enabling the model to amend flawed reasoning processes and arrive at the correct answer by comparing candidates. EVLM(khalid2024evlm), on the other hand, introduces a multi-step reflective reasoning framework that includes initial reasoning, intermediate output generation, and reflective refinement, allowing the model to internalize and generalize transformation principles through iterative reasoning. The essence of such methods lies in transforming a one-shot generation process, which may get stuck in a local optimum, into an iterative optimization process where the model progressively approaches a reasonable answer through self-critique. Furthermore, to solve the lack of visual memory and multi-step visual logic modeling capabilities in GUI agent reasoning, InfiGUIAgent(liu2025infiguiagent) utilizes a cyclical process of Expectation-Reflection reasoning to enhance the GUI agent’s behavioral consistency and self-correction abilities.
@@ 3.3.4 Knowledge Enhancement
As shown in subfigure (d) of Figure 5, some recent work enhances the reasoning capabilities of LLMs by introducing external knowledge and integrating it with the input. Such methods aim to overcome the static nature (inability to be updated in a timely manner) and unreliability (the hallucination problem) of the model’s parameterized knowledge, enabling the model to utilize more accurate and timely external facts to substantiate its reasoning process. RAGAR(khaliq2024ragar) and RMR(tan2024retrieval) utilize the Retrieval-Augmented Generation (RAG) method to incorporate domain-specific knowledge during the reasoning process to enhance the model’s reasoning capabilities. The MR-MKG method(lee2024multimodal) significantly enhances the reasoning capabilities of MLLMs by introducing Multimodal Knowledge Graphs (KGs) to learn rich semantic knowledge across modalities.
In the context of MCoT reasoning, the KAM-CoT framework(mondal2024kam) integrates CoT reasoning, knowledge graphs, and multimodal information. By introducing external knowledge from knowledge graphs during the reasoning process, it achieves a deeper contextual understanding, thereby reducing model hallucinations and improving the quality of the answers. Furthermore, Liu et al.(liu2023retrieval) proposed a retrieval-augmented CoT reasoning method that utilizes cross-modal and intra-modal similarities to dynamically select demonstration examples, which are then structured into the CoT paradigm and included as part of the input.
@@ 3.3.5 Agent Assistance
In multimodal scenarios, Elhenawy et al.(elhenawy2024visual) proposed a multi-agent method to solve the Traveling Salesperson Problem (TSP) and the multi-TSP (mTSP). Their proposed ”Multi-Agent 1” includes an initializer, a critic, and a scorer agent, while ”Multi-Agent 2” comprises an initializer and a critic agent; both multi-agent frameworks significantly improved the quality of problem-solving. Insight-V(dong2025insight) employs a multi-agent system that breaks down the solution process into two steps: reasoning (generating a detailed reasoning process) and summarizing (identifying key information and selectively answering). This division of labor reduces the cognitive load on a single agent, allowing it to focus on its core task, thereby enhancing overall efficiency and output quality. Additionally, the works of MultiMedRes(gu2024inquire), InfiGUIAgent(liu2025infiguiagent), and Zhai et al.(zhai2024fine) utilize a single agent to enhance the model’s visual reasoning capabilities.
@@ 3.4.1 Information Representation
To address this problem, CoT makes the visual perception process explicit by compelling the model to generate intermediate text, which translates key objects, relationships, and other information from the image into linguistic descriptions. This process not only creates a shared semantic space for subsequent logical operations but, more importantly, it provides anchor points for each step of the reasoning to trace back and verify visual evidence. This achieves a continuous grounding of linguistic symbols to visual facts, thereby reducing the representation gap. This step-by-step verification mechanism establishes a soft constraint within the model’s feature space, promoting the creation of a precise mapping between low-level perceptual features and high-level semantic representations. Specifically, methods such as ICoT(gao2025interleaved) innovatively incorporate the image itself as part of a reasoning step, while CoS(liu2024chain) guides the model to focus on key regions within the image, and the IoT method(zhou2024image) even enables the model to actively plan the extraction of visual information. The successful application of these methods validates the effectiveness of using explicit intermediate steps to enhance cross-modal understanding and alignment, which in turn significantly improves the factual consistency of the model’s answers and effectively suppresses the generation of model hallucinations.
@@ 3.4.2 Structured Reasoning
CoT reshapes the model’s reasoning paradigm. Its primary mechanism is task decomposition, which breaks down a complex, one-step reasoning task into a series of simpler, more controllable logical sub-steps. This approach significantly reduces the model’s cognitive load at any single decision point and effectively suppresses the compounding errors that readily occur in autoregressive generation. In practice, this decomposition process prompts the model to shift from the fast, intuitive ”System 1” mode of thinking to a slower, more deliberate, and logically rigorous ”System 2” mode. As demonstrated by Kojima et al..(kojima2022large), even a simple meta-instruction can effectively trigger this mode switch, thereby activating the latent reasoning capabilities the model has already acquired during pre-training.
Furthermore, this explicit, modular reasoning structure provides the necessary foundation for more advanced computational paradigms. It opens up a systematically explorable solution space for advanced topological structures like Tree of Thoughts(long2023large; yao2023tree) and Graph of Thoughts(besta2024graph; lei2023boosting), as well as for search strategies such as Beam Search(xu2024llava) and Monte Carlo Tree Search(dong2024progressive; yao2024mulberry). It also provides the basis for collaborative, division-of-labor reasoning in agent-assisted frameworks like Insight-V(dong2025insight). Concurrently, at each step of the reasoning, methods such as KAM-CoT(mondal2024kam) can also precisely inject external knowledge through knowledge augmentation, further enriching and calibrating the reasoning process.

### v2

@@ 3.3.1 CoT Prompting
Compared to directly generating an answer, CoT prompting, through few-shot example learning or zero-shot prompts, induces the model to spontaneously form a reasoning chain, improving its performance on multi-step reasoning tasks. Wei et al.(Wei et al., 2022) first introduced CoT prompting method in LLMs, improving model performance on arithmetic, commonsense, and logical reasoning tasks. Kojima et al.(Kojima et al., 2022) further proposed the Zero-shot CoT prompting method, which only requires adding the cognitive trigger instruction ”Let’s think step by step” to the prompt to activate the model’s internal serialization and logical inference pathways, successfully guiding its reasoning pattern.
In multimodal scenarios, the core challenge for CoT prompting lies in providing a guiding framework to synergistically process and transform information from different modalities to bridge the representation gap. The BBA method(Zhao et al., 2024a) guides a vision-language model to create separate reasoning chains for visual representations and Domain-Specific Language (DSL) representations, and then integrates the information from different modalities by aligning inconsistencies; this decoupling-alignment strategy reduces cognitive load and problem complexity. CCoT(Mitra et al., 2024) first utilizes a Multimodal Large Language Model to generate a scene graph, which is then embedded into the prompt to generate an answer. The IoT prompting method by Zhou et al.(Zhou et al., 2024a) guides the model to automatically design visual information extraction operations based on the image and question, acquiring visual auxiliary information for each reasoning step. This transforms the model from a passive receiver of information into an active agent that acquires information, making the reasoning results more credible. Furthermore, prompting methods such as DDCoT(Zheng et al., 2023), MI-CoT(Anand et al., 2024b), and ICoT(Gao et al., 2025) incorporate visual cues in the prompts, compelling the model to trace back and verify visual evidence at each step of language generation. This achieves a continuous grounding of linguistic symbols with visual information, effectively suppressing model hallucinations.s.
@@ 3.3.2 Search Strategies
Autoregressive generation models, during inference, follow a greedy strategy, meaning they select the locally optimal next token at each step. This decision-making mechanism is highly susceptible to compounding errors, where minor early deviations are continuously amplified in subsequent steps, ultimately deviating from the correct reasoning trajectory. Search strategies in CoT reasoning aim to systematically overcome this limitation. The core idea is to expand the single, fragile reasoning path into an explorable solution space at the cost of increased computation, thereby finding a more optimal reasoning path globally and improving the correctness and stability of the output. As shown in subfigure (b) of Figure 5, compared to single-path reasoning, search strategies enable the model to more comprehensively explore plausible reasoning chains by evaluating and selecting from multiple branches, enhancing its robustness and generalization capabilities in multi-step reasoning tasks. The search process for MCoT is more complex than for its textual counterpart, as it requires joint decision-making across dimensions such as modal attention, cross-modal alignment, and perceptual understanding, resulting in a higher-dimensional search space. Common search strategies include Beam Search, Monte Carlo Tree Search (MCTS), and Best-of-N (BoN) search(Amini et al., 2024; Wang et al., 2024f).
Beam Search: As a widely used heuristic search algorithm, its core idea is to retain a fixed number (known as the ”beam width”) of the most probable candidate sequences at each step of the generation process and to expand upon them in subsequent steps. Theoretically, it is a practical compromise between greedy search and exhaustive search, mitigating the fragility of a single path through limited parallel exploration. However, its heuristic assumption may lead to missing the global optimum due to short-sightedness. In the multimodal context, to balance the granularity and complexity of the search, Xu et al.(Xu et al., 2024a) proposed a stage-level beam search method, which executes the search by leveraging the phased output characteristic of LLaVA-CoT, with its performance improving as the beam width increases.
@@
To address the reliability issue of reasoning path expansion in MCoT, AR-MCTS(Dong et al., 2024) integrates MCTS with an active retrieval mechanism to achieve automatic generation of step-by-step annotations. Its innovation lies in combining internal contemplation (MCTS) with external knowledge seeking (active retrieval), ensuring the search process is both logically coherent and factually grounded, thus enhancing the reliability of the exploration. Similarly, to tackle the search effectiveness and efficiency problems in multimodal search methods, Yao et al.(Yao et al., 2024) proposed CoMCTS, which introduces collective learning into the tree search. It utilizes the collective knowledge of multiple models to collaboratively hypothesize, search, and identify effective reasoning paths through iterative operations, thereby smoothing out the cognitive biases of a single model and improving search coverage and the probability of finding the optimal solution.
To balance the performance and efficiency of AR-MCTS and CoMCTS, Wu et al.(Wu et al., 2025) proposed AStar, a multimodal automatic structured thought paradigm based on MCTS. Its contribution is to elevate the search object from low-dimensional token sequences to the level of high-dimensional, pre-defined cognitive actions, allowing for the derivation of high-level cognitive reasoning patterns with greater efficiency. Furthermore, Sun et al.(Sun et al., 2025c) proposed a simulation-based tree search algorithm. Considering that MLLMs are affected by hallucinations and struggle to produce reliable rewards, this method employs a simulation-based reward mechanism to replace the traditional one in MCTS, providing a more reliable signal for the value assessment (backpropagation step) of MCTS and enhancing the robustness of the search process.
Best-of-N Search: The basic idea of Best-of-N is that the model generates N possible outputs, and then the optimal solution is selected from them based on a preset evaluation criterion. This method increases the likelihood of obtaining a high-quality output by increasing the number of candidate solutions, making it suitable for tasks with high requirements for the quality of the generated results. The SPECULATIVE REJECTION method(Sun et al., 2024a) improves upon Best-of-N by halting the generation of low-scoring candidate answers, thereby significantly reducing the model’s demand for computational resources. The TreeBoN method(Qiu et al., 2024) merges a speculative tree search strategy with the Best-of-N method, achieving a balance between high output quality and efficient inference by iteratively expanding branches and pruning low-quality answers. Lin et al.(Lin et al., 2025c) systematically studied inference-time scaling methods like Best-of-N, demonstrating that the performance of MCoT reasoning consistently surpasses that of text-only reasoning.
@@
For challenging multimodal mathematical reasoning tasks, the AtomThink framework(Xiang et al., 2024a) combines four search strategies and a policy reward model. It divides the search into path search (using majority voting and Best-of-N to aggregate scores and find the optimal solution) and step-wise search (gradually expanding the atomic action sampling space from an initial path, using beam search and greedy strategies to prune low-quality paths). To generate model answers with better visual understanding, Wang et al.(Wang et al., 2024g) proposed VisVM to serve as a signal to guide the model’s search during inference. The innovation of VisVM lies in introducing a more forward-looking guidance signal, providing long-term value to avoid local optima.
Additionally, to address the problems of granularity mismatch and interference from irrelevant information in in-context learning, Zhang et al.(Zhang et al., 2025a) proposed BoostStep, which improves reasoning accuracy through step-level alignment in in-context learning and can be seamlessly integrated with CoT and tree search algorithms. To mimic the interleaved nature of the human visual-language reasoning process, Wang et al.(Wang et al., 2025h) proposed the VisuoThink reasoning framework, which dynamically integrates visual and textual information. Its core is a predictive rolling search mechanism that fuses visual-textual thought, enabling timely corrections by predicting intermediate state results, thereby enhancing the model’s reasoning performance.
@@ 3.3.3 Self Refinement
In recent years, a significant body of work(Dhuliawala et al., 2023; Gou et al., 2023; Paul et al., 2023; Madaan et al., 2023; Li et al., 2023b; Shinn et al., 2023) has demonstrated that large models possess self-correction capabilities; that is, they can improve the quality of their answers through reflection. This paradigm transforms the model’s generation process from a linear ”think-output” pattern to an iterative ”think-output-review-refine” loop, optimizing the reasoning path through an internal feedback mechanism without updating the model’s parameters. As illustrated in subfigure (c) of Figure 5, after generating an initial response, the model reflects upon the original input and its initial answer to generate feedback. Finally, the model refines the initial response based on this feedback. Generally, self-optimization can rely on the model’s internal prompting and external information.
To address the issue that MLLMs struggle to learn effectively from erroneous reasoning, R3V(Cheng et al., 2024) introduces self-correction and self-selection, enabling the model to amend flawed reasoning processes and arrive at the correct answer by comparing candidates. EVLM(Khalid et al., 2024), on the other hand, introduces a multi-step reflective reasoning framework that includes initial reasoning, intermediate output generation, and reflective refinement, allowing the model to internalize and generalize transformation principles through iterative reasoning. The essence of such methods lies in transforming a one-shot generation process, which may get stuck in a local optimum, into an iterative optimization process where the model progressively approaches a reasonable answer through self-critique. Furthermore, to solve the lack of visual memory and multi-step visual logic modeling capabilities in GUI agent reasoning, InfiGUIAgent(Liu et al., 2025c) utilizes a cyclical process of Expectation-Reflection reasoning to enhance the GUI agent’s behavioral consistency and self-correction abilities.
@@ 3.3.4 Knowledge Enhancement
As shown in subfigure (d) of Figure 5, some recent work enhances the reasoning capabilities of LLMs by introducing external knowledge and integrating it with the input. Such methods aim to overcome the static nature (inability to be updated in a timely manner) and unreliability (the hallucination problem) of the model’s parameterized knowledge, enabling the model to utilize more accurate and timely external facts to substantiate its reasoning process. RAGAR(Khaliq et al., 2024) and RMR(Tan et al., 2024b) utilize the Retrieval-Augmented Generation (RAG) method to incorporate domain-specific knowledge during the reasoning process to enhance the model’s reasoning capabilities. The MR-MKG method(Lee et al., 2024) significantly enhances the reasoning capabilities of MLLMs by introducing Multimodal Knowledge Graphs (KGs) to learn rich semantic knowledge across modalities.
In the context of MCoT reasoning, the KAM-CoT framework(Mondal et al., 2024) integrates CoT reasoning, knowledge graphs, and multimodal information. By introducing external knowledge from knowledge graphs during the reasoning process, it achieves a deeper contextual understanding, thereby reducing model hallucinations and improving the quality of the answers. Furthermore, Liu et al.(Liu et al., 2023a) proposed a retrieval-augmented CoT reasoning method that utilizes cross-modal and intra-modal similarities to dynamically select demonstration examples, which are then structured into the CoT paradigm and included as part of the input.
@@ 3.3.5 Agent Assistance
In multimodal scenarios, Elhenawy et al.(Elhenawy et al., 2024) proposed a multi-agent method to solve the Traveling Salesperson Problem (TSP) and the multi-TSP (mTSP). Their proposed ”Multi-Agent 1” includes an initializer, a critic, and a scorer agent, while ”Multi-Agent 2” comprises an initializer and a critic agent; both multi-agent frameworks significantly improved the quality of problem-solving. Insight-V(Dong et al., 2025) employs a multi-agent system that breaks down the solution process into two steps: reasoning (generating a detailed reasoning process) and summarizing (identifying key information and selectively answering). This division of labor reduces the cognitive load on a single agent, allowing it to focus on its core task, thereby enhancing overall efficiency and output quality. Additionally, the works of MultiMedRes(Gu et al., 2024), InfiGUIAgent(Liu et al., 2025c), and Zhai et al.(Zhai et al., 2024) utilize a single agent to enhance the model’s visual reasoning capabilities.
@@ 3.4.1 Information Representation
To address this problem, CoT makes the visual perception process explicit by compelling the model to generate intermediate text, which translates key objects, relationships, and other information from the image into linguistic descriptions. This process not only creates a shared semantic space for subsequent logical operations but, more importantly, it provides anchor points for each step of the reasoning to trace back and verify visual evidence. This achieves a continuous grounding of linguistic symbols to visual facts, thereby reducing the representation gap. This step-by-step verification mechanism establishes a soft constraint within the model’s feature space, promoting the creation of a precise mapping between low-level perceptual features and high-level semantic representations. Specifically, methods such as ICoT(Gao et al., 2025) innovatively incorporate the image itself as part of a reasoning step, while CoS(Liu et al., 2024c) guides the model to focus on key regions within the image, and the IoT method(Zhou et al., 2024a) even enables the model to actively plan the extraction of visual information. The successful application of these methods validates the effectiveness of using explicit intermediate steps to enhance cross-modal understanding and alignment, which in turn significantly improves the factual consistency of the model’s answers and effectively suppresses the generation of model hallucinations.
@@ 3.4.2 Structured Reasoning
CoT reshapes the model’s reasoning paradigm. Its primary mechanism is task decomposition, which breaks down a complex, one-step reasoning task into a series of simpler, more controllable logical sub-steps. This approach significantly reduces the model’s cognitive load at any single decision point and effectively suppresses the compounding errors that readily occur in autoregressive generation. In practice, this decomposition process prompts the model to shift from the fast, intuitive ”System 1” mode of thinking to a slower, more deliberate, and logically rigorous ”System 2” mode. As demonstrated by Kojima et al..(Kojima et al., 2022), even a simple meta-instruction can effectively trigger this mode switch, thereby activating the latent reasoning capabilities the model has already acquired during pre-training.
Furthermore, this explicit, modular reasoning structure provides the necessary foundation for more advanced computational paradigms. It opens up a systematically explorable solution space for advanced topological structures like Tree of Thoughts(Long, 2023; Yao et al., 2023b) and Graph of Thoughts(Besta et al., 2024; Lei et al., 2023), as well as for search strategies such as Beam Search(Xu et al., 2024a) and Monte Carlo Tree Search(Dong et al., 2024; Yao et al., 2024). It also provides the basis for collaborative, division-of-labor reasoning in agent-assisted frameworks like Insight-V(Dong et al., 2025). Concurrently, at each step of the reasoning, methods such as KAM-CoT(Mondal et al., 2024) can also precisely inject external knowledge through knowledge augmentation, further enriching and calibrating the reasoning process.